OpenStudy (anonymous):

a restaurant bought two 1/5 pages ads, nine 1/20 ads, and five 1/4 pages ads. what is the total amount of spaces they bought? write a number sentence.. thank you for your help.

OpenStudy (anonymous):

i know that the two 1/5 pages= 2/5

OpenStudy (anonymous):

i know that the five 1/4 pages = 1 1/4

OpenStudy (anonymous):

but is it the nine 1/20 = 9/20? and then what?

OpenStudy (anonymous):

Sum all of them.

OpenStudy (anonymous):

am i correct on the 9/20?

OpenStudy (anonymous):

is 9/20+5/4+2/5=21/10

OpenStudy (anonymous):

opps sorry 9/20 plus 2/5 = 17/20

OpenStudy (anonymous):

17/20 plus 5/4 = 42/20?? am i doing this right?

OpenStudy (anonymous):

yes.. and i simplify 42/20 .. i think i got it

OpenStudy (anonymous):

Yeah is right 17/20+5/4=(17*4+20*5)/(20*4)=(68+100)/80=(168)/80 Then take the half of both 84/40 half again 42/20 Half again 21/10

OpenStudy (anonymous):

thank u